# FIRM: Federated In-client Regularized Multi-objective Alignment for Large Language Models

### v1

@@ A.1 Experimental Setup
We conduct experiments on two publicly available Large Language Models (LLMs): meta-llama/Llama-3.2-1B-Instruct and TinyLlama/TinyLlama-1.1B-Chat-v1.0. The choice of these models is motivated by our federated learning setting, as such models are more suitable for deployment on resource-constrained edge devices. To ensure computational efficiency, we employ Parameter-Efficient Fine-Tuning (PEFT) using Low-Rank Adaptation (LoRA) (hu2021lora). For generating responses, we use prompts from the Anthropic Helpfulness and Harmlessness (HH) dataset (bai2022training).
@@
Our implementation is built using the TRL library (vonwerra2022trl). For the PPO algorithm, we set the actor and critic learning rates to $6\times 10^{-5}$ and $1\times 10^{-4}$, respectively. Each client performs 3 local PPO epochs per round, using a batch size of 16 and a minibatch size of 8. We use an adaptive KL controller with a target KL of $0.03$ to stabilize training. For efficient fine-tuning, LoRA is applied to all projection layers (q_proj, k_proj, v_proj, o_proj) with a rank of $r=16$. The regularization parameter for our MGDA-based solver was set to $\beta_{\text{mgda}}=0.01$.
@@ Appendix B Related Work
In centralized settings, research has focused on adapting reinforcement learning (RL) to handle multiple, often conflicting, reward signals. A primary challenge is resolving conflicting gradients, where an update improving one objective may degrade another. Gradient-based methods such as GAPO (li2025gradient) explicitly apply MGDA to find a common descent direction that guarantees no single objective is worsened. An alternative architectural philosophy is taken by EMORL, which avoids the gradient conflict problem altogether by training $M$ separate, specialized models—one per objective—and combining them at inference time through hidden-state aggregation kong2025emorl. While this approach sidesteps multi-objective optimization during training, it introduces substantial memory and inference overhead compared to single-model methods such as GAPO and our own. Beyond alignment-specific work, foundational research in multi-objective RL has established finite-time convergence guarantees for actor-critic methods using MGDA in the single-agent setting (zhou2024finite), reinforcing our choice of MGDA as a theoretically principled optimization technique.
A practical enabler for many recent alignment approaches, including our own, is Low-Rank Adaptation (LoRA) (hu2021lora). LoRA makes parameter-efficient fine-tuning possible by injecting low-rank matrices into transformer layers, drastically reducing the number of training parameters. This not only lowers the computational and memory footprint in centralized training but also makes federated fine-tuning of large models feasible, since clients can update and transmit only adapter weights rather than full model parameters. Our method builds on this paradigm to ensure scalability in LLM alignment experiments.
@@
Adapting alignment to a federated setting (sometimes called Fed-RLHF) has led to several architectural approaches. It is important to distinguish our work from frameworks focused on pluralistic alignment (chen2024integration; srewa2025pluralllm), which aim to aggregate diverse and subjective user preferences into a single model. Pluralistic and personalized frameworks, such as FedBiscuit (wu2024towards), directly address preference heterogeneity across clients. However, in these methods, the final stage of aligning the language model’s policy is still performed centrally. By contrast, our work targets a fully federated regime in which clients collaboratively align the policy itself, rather than only sharing preference signals.
@@
The most direct competitors to our work are algorithms that explicitly combine FL with multi-objective optimization. Most existing FMOO frameworks follow a “Client-Computes, Server-Resolves” pattern, where the server is responsible for resolving objective conflicts. Foundational work by yang2023federated exemplifies this approach: clients transmit local gradients and the server solves the multi-objective problem. A naive adaptation of this strategy to our alignment problem would require each client to send $M$ gradients, leading to a communication cost of $O(Md)$. FedCMOO (askin2024federated) reduces this burden with compression, but its randomized SVD step imposes heavy computational overhead on clients, and reconstructing gradients at the server introduces an additional source of error.
@@
Federated reinforcement learning has been studied with a focus on efficiency and convergence in single-objective settings (zhang2024finitetime; woo2024federated; khodadadian2022federated). In parallel, RL theory has established provably efficient algorithms for exploration and value estimation in centralized settings (zanette2020learning; jin2020provably; roknilamouki2025provably).
@@ Appendix D Helpful Lemmas
We first rewrite Lemma 2 from zhou2024finite:
@@
Lemma 3 (Critic Convergence Theorem 1 from xu2020improving).
@@
Note that the proof of this Lemma without the projection step in Algorithm 3 is provided in Theorem 1 of xu2020improving. However, since based on Lemma 8 in zhou2024finite we have $\|w_{t}^{*,j,c}\|_{2}\leq R_{w}$, projection step in our Alrogithm is non-expansive and the proof of Theorem 1 in xu2020improving holds for our case as well. $\square$
@@ F.1.1 Bounding the Gradient Estimation Error ($\mathcal{T}_{1,1}^{\text{gradient-error}}$)
First inequality is same as before, and here we only need to prove the second inequality. From Appendix D in zhou2024finite, we have the per-client variance bound:
@@ F.1.4 Boudning the term $\mathcal{T}_{1,2}$
To bound this term, we can use the proof of Lemma 9. First of all, by applying Proposition 1 from xu2020improving we have:
@@ G.1 Proof of Lemma 8
Now, let $\lambda^{\prime*}$ be the minizer of function $f_{G^{\prime}}(\lambda)$ and recall that $\lambda^{*}$ is the minimizer of $f_{G}(\lambda)$. Now, consdierng Equation(38) and Equation(40), we can use Proposition 4.32 from bonnans2013perturbation to get the following:
@@ G.2 Proof of Lemma 9
Note that the terms $\sqrt{E\Big[\|w_{t}^{*,c^{\prime}}-w_{t}^{c^{\prime}}\|_{2}\mid\mathcal{F}_{t}\Big]}$ and $\sqrt{E\Big[\|w_{t}^{*,c}-w_{t}^{c}\|_{2}\mid\mathcal{F}_{t}\Big]}$ represents the critic estimation error of each client and can be bounded by proper batch size of the critics. Now, in order to bound the term $E\Big[\|\nabla_{\theta}J(\theta_{t}^{c})-\nabla_{\theta}J(\theta_{t}^{c^{\prime}})\|_{2}\mid\mathcal{F}_{t}\Big]$ in Equation(48) we apply Proposition 1 from xu2020improving and will get the following:

### v2

@@ A.1 Experimental Setup
We conduct experiments on two publicly available Large Language Models (LLMs): meta-llama/Llama-3.2-1B-Instruct and TinyLlama/TinyLlama-1.1B-Chat-v1.0. The choice of these models is motivated by our federated learning setting, as such models are more suitable for deployment on resource-constrained edge devices. To ensure computational efficiency, we employ Parameter-Efficient Fine-Tuning (PEFT) using Low-Rank Adaptation (LoRA) (Hu et al., 2021). For generating responses, we use prompts from the Anthropic Helpfulness and Harmlessness (HH) dataset (Bai et al., 2022).
@@
Our implementation is built using the TRL library (von Werra et al., 2020). For the PPO algorithm, we set the actor and critic learning rates to $6\times 10^{-5}$ and $1\times 10^{-4}$, respectively. Each client performs 3 local PPO epochs per round, using a batch size of 16 and a minibatch size of 8. We use an adaptive KL controller with a target KL of $0.03$ to stabilize training. For efficient fine-tuning, LoRA is applied to all projection layers (q_proj, k_proj, v_proj, o_proj) with a rank of $r=16$. The regularization parameter for our MGDA-based solver was set to $\beta_{\text{mgda}}=0.01$.
@@ Appendix B Related Work
In centralized settings, research has focused on adapting reinforcement learning (RL) to handle multiple, often conflicting, reward signals. A primary challenge is resolving conflicting gradients, where an update improving one objective may degrade another. Gradient-based methods such as GAPO (Li et al., 2025) explicitly apply MGDA to find a common descent direction that guarantees no single objective is worsened. An alternative architectural philosophy is taken by EMORL, which avoids the gradient conflict problem altogether by training $M$ separate, specialized models—one per objective—and combining them at inference time through hidden-state aggregation Kong et al. (2025). While this approach sidesteps multi-objective optimization during training, it introduces substantial memory and inference overhead compared to single-model methods such as GAPO and our own. Beyond alignment-specific work, foundational research in multi-objective RL has established finite-time convergence guarantees for actor-critic methods using MGDA in the single-agent setting (Zhou et al., 2024), reinforcing our choice of MGDA as a theoretically principled optimization technique.
A practical enabler for many recent alignment approaches, including our own, is Low-Rank Adaptation (LoRA) (Hu et al., 2021). LoRA makes parameter-efficient fine-tuning possible by injecting low-rank matrices into transformer layers, drastically reducing the number of training parameters. This not only lowers the computational and memory footprint in centralized training but also makes federated fine-tuning of large models feasible, since clients can update and transmit only adapter weights rather than full model parameters. Our method builds on this paradigm to ensure scalability in LLM alignment experiments.
@@
Adapting alignment to a federated setting (sometimes called Fed-RLHF) has led to several architectural approaches. It is important to distinguish our work from frameworks focused on pluralistic alignment (Chen et al., 2024; Srewa et al., 2025), which aim to aggregate diverse and subjective user preferences into a single model. Pluralistic and personalized frameworks, such as FedBiscuit (Wu et al., 2024), directly address preference heterogeneity across clients. However, in these methods, the final stage of aligning the language model’s policy is still performed centrally. By contrast, our work targets a fully federated regime in which clients collaboratively align the policy itself, rather than only sharing preference signals.
@@
The most direct competitors to our work are algorithms that explicitly combine FL with multi-objective optimization. Most existing FMOO frameworks follow a “Client-Computes, Server-Resolves” pattern, where the server is responsible for resolving objective conflicts. Foundational work by Yang et al. (2023) exemplifies this approach: clients transmit local gradients and the server solves the multi-objective problem. A naive adaptation of this strategy to our alignment problem would require each client to send $M$ gradients, leading to a communication cost of $O(Md)$. FedCMOO (Askin et al., 2024) reduces this burden with compression, but its randomized SVD step imposes heavy computational overhead on clients, and reconstructing gradients at the server introduces an additional source of error.
@@
Federated reinforcement learning has been studied with a focus on efficiency and convergence in single-objective settings (Zhang et al., 2024; Woo et al., 2024; Khodadadian et al., 2022). In parallel, RL theory has established provably efficient algorithms for exploration and value estimation in centralized settings (Zanette et al., 2020; Jin et al., 2020; Roknilamouki et al., 2025).
@@ Appendix D Helpful Lemmas
We first rewrite Lemma 2 from Zhou et al. (2024):
@@
Lemma 3 (Critic Convergence Theorem 1 from Xu et al. (2020a)).
@@
Note that the proof of this Lemma without the projection step in Algorithm 3 is provided in Theorem 1 of Xu et al. (2020a). However, since based on Lemma 8 in Zhou et al. (2024) we have $\|w_{t}^{*,j,c}\|_{2}\leq R_{w}$, projection step in our Alrogithm is non-expansive and the proof of Theorem 1 in Xu et al. (2020a) holds for our case as well. $\square$
@@ F.1.1 Bounding the Gradient Estimation Error ($\mathcal{T}_{1,1}^{\text{gradient-error}}$)
First inequality is same as before, and here we only need to prove the second inequality. From Appendix D in Zhou et al. (2024), we have the per-client variance bound:
@@ F.1.4 Boudning the term $\mathcal{T}_{1,2}$
To bound this term, we can use the proof of Lemma 9. First of all, by applying Proposition 1 from Xu et al. (2020a) we have:
@@ G.1 Proof of Lemma 8
Now, let $\lambda^{\prime*}$ be the minizer of function $f_{G^{\prime}}(\lambda)$ and recall that $\lambda^{*}$ is the minimizer of $f_{G}(\lambda)$. Now, consdierng Equation(38) and Equation(40), we can use Proposition 4.32 from Bonnans and Shapiro (2013) to get the following:
@@ G.2 Proof of Lemma 9
Note that the terms $\sqrt{E\Big[\|w_{t}^{*,c^{\prime}}-w_{t}^{c^{\prime}}\|_{2}\mid\mathcal{F}_{t}\Big]}$ and $\sqrt{E\Big[\|w_{t}^{*,c}-w_{t}^{c}\|_{2}\mid\mathcal{F}_{t}\Big]}$ represents the critic estimation error of each client and can be bounded by proper batch size of the critics. Now, in order to bound the term $E\Big[\|\nabla_{\theta}J(\theta_{t}^{c})-\nabla_{\theta}J(\theta_{t}^{c^{\prime}})\|_{2}\mid\mathcal{F}_{t}\Big]$ in Equation(48) we apply Proposition 1 from Xu et al. (2020a) and will get the following: